# Improved Stochastic Optimization of LogSumExp

### v1

@@ 1 Introduction
LogSumExp function appears commonly in optimization objectives, e.g., multiclass classification with softmax probabilities (Bishop & Nasrabadi, 2006), semi-dual formulation of entropy-regularized optimal transport (OT) (Peyré & Cuturi, 2019; Genevay et al., 2016), minimax problems (Pee & Royset, 2011), distributionally robust optimization (DRO) (Hu & Hong, ; Ben-Tal et al., 2013; Kuhn et al., 2024), maximum likelihood estimation (MLE) for exponential families and graphical models (Wainwright et al., 2008), variational Bayesian methods (Khan & Nielsen, 2018; Khan & Rue, 2023), information geometry (Amari & Nagaoka, 2000), KL-regularized Markov decision processes (Tiapkin et al., 2024). Such optimization problems are characterized by two challenges. First, the decision variable $\varphi$ (or a parameter $\theta$ defining $\varphi$) often has large or infinite dimension. Second, the support of the measure $\mu$ can be large or even infinite. The first challenge is usually addressed by the use of first-order methods, especially stochastic gradient descent (SGD), with cheap iterations. Unfortunately, there is no universal way to tackle the second challenge. If the number of exponential terms under the logarithm is large or infinite, i.e., the Monte Carlo estimate $\log\sum_{i=1}^{N}e^{\varphi(x)}$ for a large $N$, the gradient computation requires differentiating each term, and, to the best of our knowledge, no cheap unbiased stochastic gradient have been proposed.
@@
In the current work, we propose a general-purpose approach to tackle both mentioned challenges. To that end, we use a SoftPlus approximation of $F(\varphi;\mu)$ that allows using stochastic gradient methods while remaining close to the original objective. We start with a variational formulation analogous to the one in the Gibbs principle, but with the KL-divergence replaced with another $f$-divergence – the safe KL divergence. The corresponding variational problem can be of interest itself, as it possesses some properties which can be beneficial compared to the KL penalty — e.g., uniform density bound. Moreover, it can be also viewed as an approximation of a conditional value at risk functional (CVaR). In fact, the same functional (with different parameters) appeared in Soma & Yoshida (2020) in the context of smooth CVaR approximation. Thus, we demonstrate that it generates a family of problems including CVaR and LogSumExp minimization as limit cases.
@@
Related works. We are not aware of any universal way to efficiently treat optimization objectives involving the LogSumExp functional (1), especially in infinite-dimensional settings. This functional appeared in different applications and was treated on a case-by-case basis. Bouchard (2007) studies three upper bounds on LogSumExp for approximate Bayesian inference. One of them is a particular case of the approximation proposed in the present work. Titsias (2016) constructs a bound on softmax probabilities and shows that it leads to a bound on LogSumExp in the context of multiclass classification. Nielsen & Sun (2016) approximate LogSumExp in the context of estimating divergences between mixture models. Their approximation combines LogSumExp bounds based on $\min$ and $\max$. Hu & Hong and, susequently, Levy et al. (2020) study DRO problems with $f$-divergences. They propose a batch-based approximation. When the ambiguity set is the unit simplex, and KL divergence penalty is used, the original objective is the LogSumExp of the losses over the entire dataset, while the approximation replaces it with the average of LogSumExp terms computed on individual batches. This approximation introduces bias, which can only be reduced by using large batch sizes.
@@ 2 SoftPlus approximation of log-partition function
$F(\varphi;\mu)=\sup\left\{\int\varphi(x)\,\mathrm{d}\nu(x)-D_{KL}(\nu,\mu):\nu\in\mathcal{P}(\mathcal{X}),\,\int|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}$ (2)
with the minimum attained at the Gibbs measure $d\nu^{*}(x)=e^{\varphi(x)-F(\varphi;\mu)}d\mu(x),$ once $F(\varphi;\mu)<\infty$ (Polyanskiy & Wu, 2025, Proposition 4.7).
@@
$D_{\rho}(\nu,\mu)\coloneqq\begin{cases}\int f_{\rho}\left(\frac{d\nu}{d\mu}(x)\right)\,\mathrm{d}\mu(x),&\nu\ll\mu,\\ +\infty,&\text{otherwise}.\end{cases}$ (4)
@@
$F_{\rho}(\varphi;\mu)\coloneqq\sup\left\{\int\varphi(x)\,\mathrm{d}\nu(x)-D_{\rho}(\nu,\mu):\right$.
$\left.\nu\in\mathcal{P}(\mathcal{X}),\,\int|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}.$ (5)
(i.e., $F_{\rho}(\cdot;\mu)$ is the convex conjugate of $D_{\rho}(\cdot,\mu)$). Note that the last term in $f_{\rho}$ prevents the density $\frac{d\nu}{d\mu}$ from being too large. In particular, it can not be greater than $\frac{1}{\rho}$. This can make the safe KL divergence a reasonable choice for unbalanced OT or DRO, as it imposes a hard constraint on the reweighting unlike the standard $D_{KL}$. Moreover, it can also be used instead of the entropy penalization in regularized OT (cf. capacity constrained transport in (Benamou et al., 2015, section 5.2)).
Again, by the convex duality and the variational principle (see Birrell et al., 2022, Theorem 6), we state the following properties.
@@
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha+\int f_{\rho}^{*}(\varphi(x)-\alpha)\,\mathrm{d}\mu(x).$
@@
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\int\log\left(1+\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x).$ (6)
@@
$\int\frac{e^{\varphi(x)-\alpha^{*}}}{1+\rho e^{\varphi(x)-\alpha^{*}}}\,\mathrm{d}\mu(x)=1,$ (7)
@@
For $\rho=1$ our approximation coincides with Bouchard’s bound for LogSumExp (Bouchard, 2007).
@@ 2.1 Links to CVaR
where $Q_{1-\rho}$ is the $(1-\rho)$-quantile of $\varphi(X)$, $X\sim\mu$ (Rockafellar et al., 2000). Moreover, by Theorem 1 in Rockafellar et al. (2000) CVaR also has the following variational formulation:
$\operatorname{CVaR}_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha+\frac{1}{\rho}\int(\varphi(x)-\alpha)_{+}\,\mathrm{d}\mu(x).$ (9)
Remarkably, in Soma & Yoshida (2020) the authors obtained a smooth approximation to CVaR which, up to an additive constant, has the same form as $F_{\rho}$. However, they considered the approximation w.r.t. a different parameter—a ”temperature” inside SoftPlus. Finally, Levy et al. (2020) proposed another similar smoothed version of CVaR (KL-regularized CVaR) in the context of DRO. For our approximation, we obtain the following bounds.
@@ 2.2 The case of parametric models
$\min_{\theta\in\Theta,\alpha\in\mathbb{R}}G_{\rho}(\theta,\alpha)\coloneqq\alpha+\log\rho-1+\frac{1}{\rho}\int\log\left(1+e^{L(x,\theta)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
$\nabla_{\theta}G_{\rho}(\theta,\alpha)=\frac{1}{\rho}\int\sigma\left(L(x,\theta)-\alpha\right)\nabla_{\theta}L(x,\theta)\,\mathrm{d}\mu(x),$
$\partial_{\alpha}G_{\rho}(\theta,\alpha)=1-\frac{1}{\rho}\int\sigma\left(L(x,\theta)-\alpha\right)\,\mathrm{d}\mu(x).$
@@ 3 Applications
In this section we consider several particular applications involving the objective (1) and show numerically, that our general-purpose approach based on approximation (6) leads to better performance of SGD-type algorithms than the baseline algorithms designed specifically for these applications. The source code for all experiments is available at https://github.com/egorgladin/logsumexp-approx.
@@ 3.1 Continuous entropy-regularized OT
The classical optimal transport (Monge–Kantorovich) problem consists in finding a coupling of two probablility measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$ which minimizes the integral of a given measurable cost function $c\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ (e.g., a distance), i.e., $W(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z)$, where $\Pi(\mu,\nu)\subset\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of couplings (transport plans) of $\mu$ and $\nu$ (see Kantorovich, 1942; Villani, 2008; Santambrogio, 2015). For simplicity of demonstration, we assume that the measures are defined on the same space $\mathcal{X}$, but the results extend trivially to the case of two different spaces. Following Cuturi (2013), we consider entropy-regularized optimal transport (eOT) problem:
$\min_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z)+\varepsilon D_{KL}(\pi,\nu\otimes\mu)$ (11)
where $\nu\otimes\mu$ is the product measure. It is known that eOT admits the following dual and semi-dual formulations (see, e.g., Genevay et al. (2016)):
$W_{\varepsilon}(\mu,\nu)=\underbrace{\max_{u,v\in\mathcal{C}(\mathcal{X})}\iint f_{\varepsilon}(x,y,u,v)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)}_{\text{dual}}=\underbrace{\max_{v\in\mathcal{C}(\mathcal{X})}\int h_{\varepsilon}(x,v)\,\mathrm{d}\mu(x)}_{\text{semi-dual}},$
@@
$h_{\varepsilon}(x,v)$ $\coloneqq\int v(y)\mathrm{d}\nu(y)-\varepsilon\log\left(\int\exp\left(\frac{v(y)-c(x,y)}{\varepsilon}\right)\mathrm{d}\nu(y)\right)-\varepsilon,$ (13)
and $\varepsilon>0$ is the regularization coefficient. Genevay et al. (2016) consider a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ defined on $\mathcal{X}$, with a kernel $\kappa$, and apply SGD to solve the dual problem.
@@
The main drawback of this approach is the presence of the exponent in the dual objective (and consequently in the SGD updates, see Appendix B.1). Specifically, exponents are prone to floating-point exceptions (Goldberg, 1991), especially if the regularization parameter $\varepsilon$ is relatively small, which is often the case. For example, if $\varepsilon=0.01$ and $z\geq 7.1$, then $e^{z/\varepsilon}$ exceeds the representable range of a double-precision (float64) floating-point number — an overflow happens. When single precision (float32) is used, an overflow happens even for $z\geq 0.89$.
@@
$\max_{\begin{subarray}{c}v\in\mathcal{C}(\mathcal{X})\\ \alpha\in\mathbb{R}\end{subarray}}\,\iint\,$ $\tilde{h}_{\varepsilon}(x,y,v,\alpha)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)$ (14)
with $\tilde{h}_{\varepsilon}(x,y,v,\alpha)\coloneqq v(y)-\alpha-\frac{\varepsilon}{\rho}\log\bigl(1+\rho e^{(v(y)-c(x,y)-\alpha)/\varepsilon}\bigr)-\varepsilon,$ (15)
which can also be solved by SGD. Analytic form of SGD iterates for objectives (12) and (15) can be found in Appendix B.1. One can show, in the same way as in Genevay et al. (2016), that this corresponds to the regularized OT problem (11) with Safe KL divergence $D_{\rho}$ rather than the usual KL, i.e.
$\min_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z)+\varepsilon D_{\rho}(\pi,\nu\otimes\mu).$
@@
Consider a setup analogous to the one described in Section 5 of Genevay et al. (2016). Specifically, $\mu$ is a 1D Gaussian, and $\nu$ is a mixture of two Gaussians (see Appendix B for a plot of densities). Gaussian kernel $\kappa(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{\sigma^{2}}\right)$ with a bandwidth hyperparameter $\sigma^{2}>0$ is used. The regularization coefficient is set to $\varepsilon=0.01$. We consider kernel SGD applied to the dual objective (12) as a baseline approach (Genevay et al., 2016). We compare it to the proposed approach, namely, kernel SGD applied to the approximate semi-dual problem (14). For details on how the optimality gap is estimated, see Appendix B.

### v2

@@ 1 Introduction
LogSumExp function appears commonly in optimization objectives, e.g., multiclass classification with softmax probabilities (bishop2006pattern), semi-dual formulation of entropy-regularized optimal transport (OT) (peyre2018computational; genevay2016stochastic), minimax problems (pee2011solving), distributionally robust optimization (DRO) (huKullbackLeiblerDivergenceConstrained; ben-tal_robust_2013; kuhnDistributionallyRobustOptimization2024), maximum likelihood estimation (MLE) for exponential families and graphical models (wainwright2008graphical), variational Bayesian methods (khan2018fast; khanBayesianLearningRule2023), information geometry (amariMethodsInformationGeometry2000), KL-regularized Markov decision processes (tiapkin2024demonstrationregularized). These problems involve minimizing $F(\varphi;\mu)$ w.r.t. a function $\varphi$, potentially parameterized by a vector $\theta$, e.g., a vector of neural network weights. Such optimization is characterized by two challenges. First, the decision variable $\varphi$ or $\theta$ often has large or infinite dimension. Second, the support of the measure $\mu$ can also be large or infinite. The first challenge is usually addressed by the use of first-order methods like Stochastic Gradient Descent (SGD), which are suitable for high-dimensional problems due to their cheap iterations. Unfortunately, there is no universal way to tackle the second challenge. If the number of exponential terms under the logarithm is large or infinite, i.e., the Monte Carlo estimate $\log\sum_{i=1}^{N}e^{\varphi(x_{i})}$ for a large $N$, the gradient computation requires differentiating each term, and, to the best of our knowledge, no cheap unbiased stochastic gradient have been proposed.
@@
In the current work, we propose a general-purpose approach to tackle both mentioned challenges. To that end, we use a SoftPlus approximation of $F(\varphi;\mu)$ that allows using stochastic gradient methods while remaining close to the original objective. We start with a variational formulation analogous to the one in the Gibbs principle, but with the KL-divergence replaced with another $f$-divergence – the safe KL divergence. The corresponding variational problem can be of interest itself, as it possesses some properties which can be beneficial compared to the KL penalty – e.g., uniform density bound. Moreover, it can be also viewed as an approximation of a conditional value at risk functional (CVaR). In fact, the same functional (with different parameters) appeared in soma2020statistical in the context of smooth CVaR approximation. Thus, we demonstrate that it generates a family of problems including CVaR and LogSumExp minimization as limit cases.
@@
Related works. We are not aware of any universal way to efficiently treat optimization objectives involving the LogSumExp functional (1), especially in infinite-dimensional settings. This functional appeared in different applications and was treated on a case-by-case basis. bouchard2007efficient studies three upper bounds on LogSumExp for approximate Bayesian inference. One of them is a particular case of the approximation proposed in the present work. aueb2016one constructs a bound on softmax probabilities and shows that it leads to a bound on LogSumExp in the context of multiclass classification. nielsen2016guaranteed approximate LogSumExp in the context of estimating divergences between mixture models. Their approximation combines LogSumExp bounds based on $\min$ and $\max$. Tucker2017REBAR; Luo2020SUMO propose and study unbiased estimators for latent variable models based on Russian Roulette truncation. LyneGirolamiEtAl2015RussianRoulette; SpringShrivastava2017LSHPartition focus on estimating the partition function itself, and do not consider questions of optimization involving the partition function. huKullbackLeiblerDivergenceConstrained and, susequently, levy2020large study DRO problems with $f$-divergences. They propose a batch-based approximation. When the ambiguity set is the unit simplex, and KL divergence penalty is used, the original objective is the LogSumExp of the losses over the entire dataset, while the approximation replaces it with the average of LogSumExp terms computed on individual batches. This approximation introduces bias, which can only be reduced by using large batch sizes. Deterministic LogSumExp maximization and minimization were considered in Selvi2020ConvexMaximization and KanNagyRuthotto2023LSEMINK, respectively. For stabilizing numerics related to evaluation of LogSumExp function, we refer to BlanchardHighamHigham2020AccurateLogSumExp; Higham2021WhatIsLogSumExp.
@@ 2 SoftPlus approximation of log-partition function
$F(\varphi;\mu)={\sup_{\nu}}\left\{\int_{{\mathcal{X}}}\varphi(x)\,\mathrm{d}\nu(x)-D_{KL}(\nu,\mu):\nu\in\mathcal{P}(\mathcal{X}),\,\int_{{\mathcal{X}}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}$ (2)
with the maximum attained at the Gibbs measure $d\nu^{*}(x)=e^{\varphi(x)-F(\varphi;\mu)}d\mu(x),$ once $F(\varphi;\mu)<\infty$, see (gibbs1902elementary, Chapter XI, Theorem VI) or (polyanskiy2025information, Proposition 4.7) for the modern treatment.
@@
$D_{\rho}(\nu,\mu)\coloneqq\begin{cases}\int_{{\mathcal{X}}}f_{\rho}\left(\frac{d\nu}{d\mu}(x)\right)\,\mathrm{d}\mu(x),&\nu\ll\mu,\\ +\infty,&\text{otherwise}.\end{cases}$ (4)
@@
$F_{\rho}(\varphi;\mu)\coloneqq\sup_{{\nu}}\left\{\int_{{\mathcal{X}}}\varphi(x)\,\mathrm{d}\nu(x)-D_{\rho}(\nu,\mu):\right$.
$\left.\nu\in\mathcal{P}(\mathcal{X}),\,\int_{{\mathcal{X}}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}.$ (5)
(i.e., $F_{\rho}(\cdot;\mu)$ is the convex conjugate of $D_{\rho}(\cdot,\mu)$). Note that the last term in $f_{\rho}$ prevents the density $\frac{d\nu}{d\mu}$ from being too large. In particular, it can not be greater than $\frac{1}{\rho}$. This can make the safe KL divergence a reasonable choice for unbalanced OT or DRO, as it imposes a hard constraint on the reweighting unlike the standard $D_{KL}$. Moreover, it can also be used instead of the entropy penalization in regularized OT (cf. capacity constrained transport in (benamou2015iterative, section 5.2)).
Again, by the convex duality and the variational principle (see birrell2022divergence, Theorem 6), we state the following properties.
@@
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha+\int_{{\mathcal{X}}}f_{\rho}^{*}(\varphi(x)-\alpha)\,\mathrm{d}\mu(x).$
@@
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\int_{{\mathcal{X}}}\log\left(1+\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x).$ (6)
@@
$\int_{{\mathcal{X}}}\frac{e^{\varphi(x)-\alpha^{*}}}{1+\rho e^{\varphi(x)-\alpha^{*}}}\,\mathrm{d}\mu(x)=1,$ (7)
@@
For $\rho=1$ our approximation coincides with Bouchard’s bound for LogSumExp (bouchard2007efficient).
@@ 2.1 Links to CVaR
where $Q_{1-\rho}$ is the $(1-\rho)$-quantile of $\varphi(X)$, $X\sim\mu$ (rockafellar2000optimization). Moreover, by Theorem 1 in rockafellar2000optimization CVaR also has the following variational formulation:
$\operatorname{CVaR}_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha+\frac{1}{\rho}\int_{{\mathcal{X}}}(\varphi(x)-\alpha)_{+}\,\mathrm{d}\mu(x).$ (9)
Remarkably, in soma2020statistical the authors obtained a smooth approximation to CVaR which, up to an additive constant, has the same form as $F_{\rho}$. However, they considered the approximation w.r.t. a different parameter—a ”temperature” inside SoftPlus. Finally, levy2020large proposed another similar smoothed version of CVaR (KL-regularized CVaR) in the context of DRO. For our approximation, we obtain the following bounds.
@@ 2.2 The case of parametric models
$\min_{\theta\in\Theta,\alpha\in\mathbb{R}}G_{\rho}(\theta,\alpha)\coloneqq\alpha+\log\rho-1+\frac{1}{\rho}\int_{{\mathcal{X}}}\log\left(1+e^{L(x,\theta)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
$\nabla_{\theta}G_{\rho}(\theta,\alpha)=\frac{1}{\rho}\int_{{\mathcal{X}}}\sigma\left(L(x,\theta)-\alpha\right)\nabla_{\theta}L(x,\theta)\,\mathrm{d}\mu(x),$
$\partial_{\alpha}G_{\rho}(\theta,\alpha)=1-\frac{1}{\rho}\int_{{\mathcal{X}}}\sigma\left(L(x,\theta)-\alpha\right)\,\mathrm{d}\mu(x).$
@@ 3 Applications
In this section we consider several particular applications involving the objective (1) and show numerically, that our general-purpose approach based on approximation (6) leads to better performance of SGD-type algorithms than the baseline algorithms designed specifically for these applications. Source code for all experiments is available at https://github.com/egorgladin/logsumexp-approx.
@@ 3.1 Continuous entropy-regularized OT
The classical optimal transport (Monge–Kantorovich) problem consists in finding a coupling of two probablility measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$ which minimizes the integral of a given measurable cost function $c\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ (e.g., a distance), i.e., $W(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z)$, where $\Pi(\mu,\nu)\subset\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of couplings (transport plans) of $\mu$ and $\nu$ (see kantorovich1942translocation; villani2008optimal; santambrogio_optimal_2015). For simplicity of demonstration, we assume that the measures are defined on the same space $\mathcal{X}$, but the results extend trivially to the case of two different spaces. Following cuturi2013sinkhorn, we consider entropy-regularized optimal transport (eOT) problem:
$\min_{\pi\in\Pi(\nu,\mu)}\int_{{\mathcal{X}\times\mathcal{X}}}c(x,z)\,\mathrm{d}\pi(x,z)+\varepsilon D_{KL}(\pi,\nu\otimes\mu)$ (11)
where $\nu\otimes\mu$ is the product measure. It is known that eOT admits the following dual and semi-dual formulations (see, e.g., genevay2016stochastic):
$W_{\varepsilon}(\mu,\nu)=\underbrace{\max_{u,v\in\mathcal{C}(\mathcal{X})}\iint_{{\mathcal{X}\times\mathcal{X}}}f_{\varepsilon}(x,y,u,v)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)}_{\text{dual}}=\underbrace{\max_{v\in\mathcal{C}(\mathcal{X})}\int_{{\mathcal{X}}}h_{\varepsilon}(x,v)\,\mathrm{d}\mu(x)}_{\text{semi-dual}},$
@@
$h_{\varepsilon}(x,v)$ $\coloneqq\int_{{\mathcal{X}}}v(y)\mathrm{d}\nu(y)-\varepsilon\log\left(\int_{{\mathcal{X}}}\exp\left(\frac{v(y)-c(x,y)}{\varepsilon}\right)\mathrm{d}\nu(y)\right)-\varepsilon,$ (13)
and $\varepsilon>0$ is the regularization coefficient. In the LSOT framework (seguy2018large), the potentials $u$ and $v$ are parameterized by neural networks and optimized via SGD. While Appendix B.1 contains a more detailed literature review, we briefly position LSOT among other solvers to motivate its selection as a baseline. LSOT offers two key advantages relevant to our goals: it is less computationally intensive than modern solvers requiring adversarial training (korotin2023neural; gushchin2023entropic; asadulaev2024neural) or iterative Langevin dynamics (mokrov2024energyguided), and it supports a general cost function—contrary to other efficient solvers like (korotin2024light) tailored to the quadratic cost. Therefore, to solve eOT with a general cost function under modest computational constraints, we adopt the LSOT framework as our primary baseline. In Appendix B.2 we compare also to (genevay2016stochastic) who use an RKHS parametrization for the potentials $u$ and $v$.
@@
The main drawback of this approach is the presence of the exponent in the dual objective (and consequently in the SGD updates). Specifically, exponents are prone to floating-point exceptions (goldberg1991every), especially if the regularization parameter $\varepsilon$ is relatively small, which is often the case. For example, if $\varepsilon=0.01$ and $z\geq 7.1$, then $e^{z/\varepsilon}$ exceeds the representable range of a double-precision (float64) floating-point number — an overflow happens. When single precision (float32) is used, an overflow happens even for $z\geq 0.89$.
@@
$\max_{v,\alpha\in\mathcal{C}(\mathcal{X})}\,\iint_{{\mathcal{X}\times\mathcal{X}}}\,$ $\tilde{h}_{\varepsilon}(x,y,v,\alpha)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)$ (14)
with $\tilde{h}_{\varepsilon}(x,y,v,\alpha)\coloneqq v(y)-\alpha(x)-\frac{\varepsilon}{\rho}\log\bigl(1+\rho e^{(v(y)-c(x,y)-\alpha(x))/\varepsilon}\bigr)-\varepsilon,$ (15)
which also admits neural network parameterization and optimization via SGD. One can show, in the same way as in genevay2016stochastic, that this corresponds to the regularized OT problem (11) with Safe KL divergence $D_{\rho}$ rather than the usual KL, i.e.
$\min_{\pi\in\Pi(\nu,\mu)}\int_{{\mathcal{X}\times\mathcal{X}}}c(x,z)\,\mathrm{d}\pi(x,z)+\varepsilon D_{\rho}(\pi,\nu\otimes\mu).$
@@
We consider the MNIST (deng2012mnist) and EMNIST-letters (cohen2017emnist) datasets as samples from the distributions $\mu$ (digits) and $\nu$ (letters). Manhattan distance $\ell_{1}$ is chosen as the cost function for computing eOT between $\mu$ and $\nu$. We parameterize the functions $u$, $v$ in LSOT and $v$, $\alpha$ in our proposed approach using a multilayer perceptron with two hidden layers (dimensions 256 and 128) and ReLU activations. The batch size is 256, and the learning rate is selected via grid search over $\{10^{-6},10^{-5},\ldots,10^{-1}\}$. The objective is evaluated on the empirical distributions of the dedicated test sets.
Figure 2 shows the performance of LSOT with the best learning rate for each regularization parameter $\varepsilon\in\{1,10^{-2},\ldots,10^{-4}\}$. It also depicts our proposed approach with the best learning rate for each $\rho\in\{10^{-1},10^{-3},10^{-5}\}$. The baseline performs adequately under strong regularization ($\varepsilon=1$). However, for weaker regularization, a learning rate of $10^{-5}$ is required to avoid numerical instability, which leads to prohibitively slow progress (red curves). Increasing the rate to $10^{-4}$ (dashed black curves) results in numerical overflow after only $\approx$5k iterations, forcing us to abort the LSOT runs at that point.
Performance of our proposed approach align with the theoretical analysis in Section 2. A large $\rho$ yields stable convergence but introduces an approximation gap, while a very small $\rho$ degrades smoothness, necessitating a smaller step size and slower training. The intermediate value $\rho=10^{-3}$ achieves the best trade-off, providing both accuracy and sufficient smoothness. In summary, our proposed approach to eOT is computationally efficient, accommodates general costs, and handles weak regularization robustly, thereby overcoming a key limitation of LSOT.